OpenStudy (darkbluechocobo):

Help Fast please with system of equations! Please! Classify the system and determine the number of solutions. (Check all that apply. ) 3x - y + 2z = 4 2x - y + 3z = 7 -9x + 3y - 6z = -12 infinitely many solutions inconsistent consistent no solution

ganeshie8 (ganeshie8):

hint : first equation and last equation are same

OpenStudy (darkbluechocobo):

How are they?

ganeshie8 (ganeshie8):

take first equation, multiply -3 both sides, wat do u get ?

OpenStudy (darkbluechocobo):

the last one :p

OpenStudy (darkbluechocobo):

so does that mean that they cancel out?

ganeshie8 (ganeshie8):

nope, when both equations are same, that means they're overlapping.... so there are infinite solutions

OpenStudy (darkbluechocobo):

:O

OpenStudy (darkbluechocobo):

Thank you :D

ganeshie8 (ganeshie8):

np :)

OpenStudy (darkbluechocobo):

Can you help me with one more?

ganeshie8 (ganeshie8):

sure :)

OpenStudy (darkbluechocobo):

Classify the system and determine the number of solutions. (Check all that apply.) x - 2y = -8 4x = 8y - 56 no solution infinitely many solutions consistent inconsisten

ganeshie8 (ganeshie8):

multiply the first equation wid 4 wat do u get ?

OpenStudy (darkbluechocobo):

4x-8y=-32

ganeshie8 (ganeshie8):

yes :) put the second equaiton also in standard form

ganeshie8 (ganeshie8):

second equation :- 4x = 8y - 56 4x - 8y = - 56

ganeshie8 (ganeshie8):

so these two equations are DIFFERENT. but clearly they have same slopes. so ?

OpenStudy (darkbluechocobo):

Basically they are parallel?

ganeshie8 (ganeshie8):

yup, they're parallel cuz they have same slope. they're NOT overlapping either

ganeshie8 (ganeshie8):

those two line are like, the train tracks.... they go forever, they never meet

ganeshie8 (ganeshie8):

so NO solution

OpenStudy (darkbluechocobo):

So that would be inconsistent no solution

ganeshie8 (ganeshie8):

Correct ! when u dont have solutions, we call it a inconsistent system

OpenStudy (darkbluechocobo):

Thank you, You have been a life saver :D

ganeshie8 (ganeshie8):

np =))

OpenStudy (darkbluechocobo):

Question Could you check this last thing for me before I submit it

ganeshie8 (ganeshie8):

sure :) you may ask any number of questions... this need not be the last thing lol ;)

OpenStudy (darkbluechocobo):

Lol :p Thank you

OpenStudy (darkbluechocobo):

I chose the third answer o.o

OpenStudy (darkbluechocobo):

because it has to be at least or equal to 90 right?

ganeshie8 (ganeshie8):

Yes ! you're right ! good job !!

OpenStudy (darkbluechocobo):

XD THank you one more sorry I just really want to do good on this test

OpenStudy (darkbluechocobo):

I chose the second one o.o

OpenStudy (darkbluechocobo):

because does it not mean when it is > or < doesn't that mean it is a dashed line instead of solid

ganeshie8 (ganeshie8):

sorry was afk,, one sec let me check :)

OpenStudy (darkbluechocobo):

lol Nevermind man :p 30 seconds before i have to submit

OpenStudy (darkbluechocobo):

I was right :p I scored a 98% :D

ganeshie8 (ganeshie8):

wow ! congrats !! but still u missed one or two it seems.... its not 100%

OpenStudy (darkbluechocobo):

I missed one :p I mixed up a definition ._.

ganeshie8 (ganeshie8):

ahh still its good :) aim for 100 next time !

OpenStudy (darkbluechocobo):

I shall :p Thank you so so so so so so much

ganeshie8 (ganeshie8):

np :D

OpenStudy (darkbluechocobo):

You really helped me out :3 I despise system of equations so much

OpenStudy (darkbluechocobo):

@ganeshie8 can we use this as help for this

ganeshie8 (ganeshie8):

yep wats the question again :)

OpenStudy (darkbluechocobo):

By the way the teacher gave my last point i missed back so... i got 100% :p

ganeshie8 (ganeshie8):

oh cool 98+1 = 100 is it :P

OpenStudy (darkbluechocobo):

xD it was worth 2 points so :p the calculation does work xD

ganeshie8 (ganeshie8):

given two equations, u knw how to write the matrix form ? Ax = b ?

OpenStudy (darkbluechocobo):

Well I have some knowledge on how to do this

OpenStudy (darkbluechocobo):

basically you have to move the variables on one side

OpenStudy (darkbluechocobo):

so \[5x + 9y = 1\] and \[4x - 7y = \]

OpenStudy (darkbluechocobo):

=2

OpenStudy (darkbluechocobo):

nevermind that is for finding the determinant .-. *sigh*

ganeshie8 (ganeshie8):

if the two equation are below :- \(a_1x + b_1y = c_1\) \(a_2x + b_2y = c_2\) then, matrix equation wud be :- \[ \left[ \begin{array}{cc} a_1 & b_1 \\ a_2 & b_2 \\ \end{array} \right] \left[ \begin{array}{cc} x \\ y \\ \end{array} \right] = \left[ \begin{array}{cc} c_1 \\ c_2 \\ \end{array} \right] \]

OpenStudy (darkbluechocobo):

ok

OpenStudy (darkbluechocobo):

so then what is the x and y ?

ganeshie8 (ganeshie8):

if the two equation are below :- \(5x+9y = 1\) \(-4x-7y = 2\) then, matrix equation wud be :- ????

ganeshie8 (ganeshie8):

you got the matrix equation part of the question ?

ganeshie8 (ganeshie8):

just from knowing the form of matrix equation, u can strike off two options : A and C

ganeshie8 (ganeshie8):

so answer is between B and D

OpenStudy (darkbluechocobo):

would it be B?

OpenStudy (darkbluechocobo):

Or is the 1 and the 2 stand or the x and y values?

ganeshie8 (ganeshie8):

obviously B is correct, but how did u get ?

OpenStudy (darkbluechocobo):

Because D has the equals as the x and y values and that is false

ganeshie8 (ganeshie8):

you simply tested whether the given x, y values are satisfying given equations or not is it ? Nice :) lets do next problem

OpenStudy (darkbluechocobo):

for the next one would it be B? Because its the right format and it doesn't use the equals as the x and y variables

ganeshie8 (ganeshie8):

B is correct ! i dint get wat do u mean by : "it doesnt use the equals as x and y variables" :|

OpenStudy (darkbluechocobo):

Well you see how the answer is 3 and 1

OpenStudy (darkbluechocobo):

c1 and c2 they are the answers to the equations that you got them from so... it obvious that it isn't the answer

ganeshie8 (ganeshie8):

okie got u :)

OpenStudy (darkbluechocobo):

the next problem is different from the other onces

ganeshie8 (ganeshie8):

you knw how to find the \(determinant\) ?

OpenStudy (darkbluechocobo):

yes

OpenStudy (darkbluechocobo):

its like cross multiplying and then subtracting

ganeshie8 (ganeshie8):

If the matrix is \[ \left[ \begin{array}{cc} a_1 & b_1 \\ a_2 & b_2 \\ \end{array} \right] \] then, its inverse is :- \[ \frac{1}{determinant} \left[ \begin{array}{cc} b_2 & -b_1 \\ -a_2 & a_1 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

memorize that for the rest of ur time in matrices chapter

OpenStudy (darkbluechocobo):

lol This is the last part of the chapter actually :p

ganeshie8 (ganeshie8):

yes, first find \(determinant\) for given matrix wat do u get ?

ganeshie8 (ganeshie8):

good you're lucky :)

OpenStudy (darkbluechocobo):

let me see .165 -.16 ??

ganeshie8 (ganeshie8):

\(determinant = \frac{1}{2}\times \frac{1}{3} - 0 \times \frac{-1}{6}\)

OpenStudy (darkbluechocobo):

thats 0

OpenStudy (darkbluechocobo):

I put them into fractions lol

OpenStudy (darkbluechocobo):

decimals *

ganeshie8 (ganeshie8):

\(determinant = \frac{1}{2}\times \frac{1}{3} - 0 \times \frac{-1}{6}\) \(=\frac{1}{6} - 0\) \(=\frac{1}{6} \)

ganeshie8 (ganeshie8):

you need to learn to be bit lazy lol dont do anything unless it is inevitable

OpenStudy (darkbluechocobo):

xD So. Keep them as fractions?

ganeshie8 (ganeshie8):

yup ! Always

OpenStudy (darkbluechocobo):

so can you do the multiplying for me i can't do fractions lol

ganeshie8 (ganeshie8):

\[Inverse ~~~ = ~~~ \frac{1}{determinant} \left[ \begin{array}{cc} b_2 & -b_1 \\ -a_2 & a_1 \\ \end{array} \right] \] \[~~~~~~~~~~~~~~~ ~~~ = ~~~ \frac{1}{\frac{1}{6}} \left[ \begin{array}{cc} \frac{1}{3} & 0 \\ \frac{1}{6} & \frac{1}{2} \\ \end{array} \right] \] \[~~~~~~~~~~~~~~~ ~~~ = ~~~ 6 \left[ \begin{array}{cc} \frac{1}{3} & 0 \\ \frac{1}{6} & \frac{1}{2} \\ \end{array} \right] \] \[~~~~~~~~~~~~~~~ ~~~ = ~~~ \left[ \begin{array}{cc} 2 & 0 \\ 1 & 3 \\ \end{array} \right] \]

ganeshie8 (ganeshie8):

see if that makes some sense

OpenStudy (darkbluechocobo):

So 6 is the determinant. You multiplied 6 by each of the numbers and got 2 0 1 3